# Linear Phase Balancing Scheme using Voltage Unbalance Sensitivities in Multi-phase Power Distribution Grids

### v1

@@ Abstract
Power distribution networks, especially in North America, are often unbalanced due to the mix of single-, two- and three-phase networks as well as due to the high penetration of single-phase devices at the distribution level such as electric vehicle (EV) chargers and single-phase solar plants. However, the network operator must adhere to the voltage unbalance levels within the limits specified by IEEE, IEC, and NEMA standards for the safety of the equipment as well as the efficiency of the network operation. Existing works have proposed active and reactive power control in the network to minimize imbalances. However, these optimization problems are highly nonlinear and nonconvex due to the inherent non-linearity of unbalanced metrics and power-flow equations. In this work, we propose a linearization approach of unbalance metrics such as voltage unbalance factors (VUF), phase voltage unbalance rate (PVUR), and line voltage unbalance rate (LVUR) using the first order Taylor’s approximation. This linearization is then applied to the phase balancing control scheme; it is formulated as a feedback approach where the linearization is updated successively after the active/reactive control setpoint has been actuated and shows improvement in voltage imbalances. We demonstrate the application of the proposed scheme on a standard IEEE benchmark test case, demonstrating its effectiveness.
Index Terms:
Phase balancing, voltage unbalance, power distribution networks, linearization.
@@ I Introduction
The power distribution systems are progressively facing increased unbalanced power injections due to the rapid growth of single-phase solar plants and residential electric vehicle chargers. This change along with the presence of inherent unbalancedness (due to mix of single and two-phase feeders) of North American grids gives rise to higher level of voltage unbalance. An increased level unbalanced in distribution systems may trigger protection devices, increase network losses, damage power equipments (e.g., distribution transformers). A network with high power unbalance results in poor utilization as one of the phases are overloaded compared to other and this may necessitate system upgrades [1]. Therefore, the system operators must satisfy the standards for regarding the unbalancedness to ensure safe and reliable operation of power distribution networks, defined by IEEE [2], IEC [3], and NEMA [4].
@@
In the existing literature, several methods are proposed to address the issue of voltage imbalance, they can be broadly categorized into three groups. The first method involves investing in new devices, such as static or dynamic VAR compensators, shunt capacitors, tap changers, voltage regulators etc. [5], which can be costly and takes time to plan and install them. The second approach is the phase switching or swapping strategy [6, 7, 8], which involves reassigning certain loads to different phases at a bus to balance loads and reduce voltage unbalance. There are also hybrid approaches that combine phase switching with the use of static VAR compensators [9]. Utilities often employ these methods manually, where maintenance crews travel to sites and physically switch loads. This process is costly, requiring extensive planning and potentially costing tens of thousands of dollars per phase swap [10]. These balancing maneuvers are typically performed periodically, as load characteristics vary with seasonal changes [11, 12, 13, 14, 15]. The third method uses active and reactive power control from existing or newly installed solar inverters and other distributed energy resources (DERs). Many studies suggest regulating the active and reactive powers from solar PV inverters [15, 16]. This strategy may require investments in the communication and situational awareness infrastructure for real-time control [17]. Most of these approaches are usually non-convex due to non-linear and non-convex nature of the unbalance metrics as well as the power flow equations. Except there are some works that have a proposed successive linearization such as in [18], convexification via linearized distflow e.g., [19, 20], etc.
@@
In this work, we propose to linearize the voltage unbalance metrics by computing the sensitivities of the voltage unbalance factors with respect to the control variables, enabling to have a linear phase balancing formulation using the first order Taylor’s approximation. These coefficients are derived for different unbalance metrics such as voltage unbalance factors (VUF), phase voltage unbalance rate (PVUR), and line voltage unbalance rate (LVUR). The work shows the accuracy of the linearization on an IEEE benchmark network, then also shows how it can be used to formulate a linearized phase balancing problem. In summary, the key contributions of the presented work are (i) analytical derivation of the sensitivities of the different voltage unbalance factors with respect to control variables such as active and reactive powers, (ii) formulating a linearized phase balancing scheme for the active and reactive power control per phase minimizing the voltage unbalance metrics.
@@
The paper is organized as follows. Section II introduces the key notations and the voltage magnitude sensitivities. Section III derives the sensitivities coefficients of different voltage unbalance metrics using the voltage magnitude sensitivities. Section IV demonstrates an application of the proposed sensitivity coefficients. Section V presents numerical validation and finally, Section VI concludes the work.
@@ II-A Notations
We consider a distribution network with network indices $i\in\mathcal{N}$, where the node indices are contained in set $\mathcal{N}=\{1,\dots,N\}$. We consider a three-phase system with phases denoted by $a,b$ and $c$. The symbols $v_{ia},v_{ib}$ and $v_{ic}$ denote the complex voltages for phase $\phi\in\Phi=\{a,b,c\}$ and the symbols $v_{iab}$, $v_{ibc}$ and $v_{ica}$ refer to line voltages. The symbols $p_{ia},p_{ib}$ and $p_{ic}$ and $q_{ia},q_{ib}$ and $q_{ic}$ denote the active and reactive powers, respectively. The symbols $|y|$ and $y^{*}$ denote magnitude and complex conjugate of $y$, respectively.
@@ II-B Voltage Sensitivities
Here $x$ can be active, reactive power or other controllable variable such as tap changers. The same coefficient can be defined for voltage magnitudes and angles. The sensitivities of magnitudes can be found as
$\frac{\partial{|v_{i\phi}|}}{\partial x_{j\phi}}=\frac{1}{|v_{i\phi}|}\Re\Big{(}v_{i\phi}^{*}\frac{\partial{v_{i\phi}}}{\partial x_{j\phi}}\Big{)}.$ (2)
These sensitivity coefficients can either computed by inverting the Jacobian of power-flow solution using the Newton-Raphson approach or by admittance matrix-based approach which relies on solving a linear set of equation [21, 22, 23] or a impedance matrix-based approach as described in [24].
@@ III Voltage Unbalance Metrics and their Sensitivities
In a multi-phase electric distribution grid, the voltage unbalance is usually referred to the cases when the three phase voltage are not symmetrical, which means that either voltage magnitudes are unequal or the voltage phase angles are not 120 apart or both of the above. This may happen due to inherent unbalancedness in the network such as the presence of single- and two-phase circuits, untransposed circuits, or the presence of the single-phase injections (load or generation). The degree of unbalance is quantified by several metrics defined by the international institutions such as IEEE[2], IEC [3] and NEMA [4]. These standards also prescribe the limits on each of these metrics for safe operation of the power distribution systems. In practice, the electric distribution system operators try their best to keep the system “balanced” within the limits defined by these standards.
In the following section, we introduce different voltage unbalance metrics and shows how their sensitivity with respect to a generic control variable can be analytically expressed.
@@ III-A1 Definition
Here, $\alpha=e^{j\frac{2\pi}{3}}$ and symbols $v_{i0},v_{i+}$ and $v_{i-}$ denote the zero, positive and negative sequence voltage respectively, for $i-$th node. The IEC standard in [3] mandates to keep the VUF value within 0.02 per unit (or 2%) for the power distribution systems.
@@ III-B1 Definition
PVUR is defined by IEEE [2], and it uses the phase to ground voltages to express the imbalance metric instead of the sequence voltages as it was for the VUF. It is defined as the ratio of (i) maximum deviation of the phase voltage magnitudes from their average and (ii) average phase voltage magnitudes. It is $\Xi_{i}^{\text{PVUR}}=$
@@
$\begin{dcases}\frac{\big{|}2|v_{ia}|-|v_{ib}|-|v_{ic}|\big{|}}{|v_{ia}|+|v_{ib}|+|v_{ic}|}\\ \qquad\quad\big{|}|v_{ia}|-{v_{p}}_{i}^{\text{avg}}\big{|}>\big{|}|v_{ib}|-{v_{p}}_{i}^{\text{avg}}\big{|}~{}\&~{}\big{|}|v_{ic}|-{v_{p}}_{i}^{\text{avg}}\big{|}\\ \frac{\big{|}2|v_{ib}|-|v_{ia}|-|v_{ic}|\big{|}}{|v_{ia}|+|v_{ib}|+|v_{ic}|}\\ \qquad\quad\big{|}|v_{ib}|-{v_{p}}_{i}^{\text{avg}}\big{|}>\big{|}|v_{ia}|-{v_{p}}_{i}^{\text{avg}}\big{|}~{}\&~{}\big{|}|v_{ic}|-{v_{p}}_{i}^{\text{avg}}\big{|}\\ \frac{\big{|}2|v_{ic}|-|v_{ia}|-|v_{ic}|\big{|}}{|v_{ia}|+|v_{ib}|+|v_{ic}|}\\ \quad\qquad\big{|}|v_{ic}|-{v_{p}}_{i}^{\text{avg}}\big{|}>\big{|}|v_{ia}|-{v_{p}}_{i}^{\text{avg}}\big{|}~{}\&~{}\big{|}|v_{ib}|-{v_{p}}_{i}^{\text{avg}}\big{|}.\\ \end{dcases}$ (8)
@@ III-B2 PVUR Sensitivities (PVURS)
Based on the three cases defined in (8), the PVUR sensitivity for each of these cases can be computed by differentiating with respect to the control variables.
The PVURS for the first case ($\big{|}|v_{ia}|-{v_{p}}_{i}^{\text{avg}}\big{|}>\big{|}|v_{ib}|-{v_{p}}_{i}^{\text{avg}}\big{|}~{}\&~{}\big{|}|v_{ic}|-{v_{p}}_{i}^{\text{avg}}\big{|}$)) can be expressed as, $\text{PVUR}_{ij\phi}=$
$&\frac{\partial(\Xi_{i}^{\text{PVUR}})}{\partial x_{j\phi}}=\frac{1}{\big{(}\sum_{i\phi}{|v_{i\phi}|}\big{)}^{2}}\Big{\{}\big{(}\sum_{i\phi}{|v_{i\phi}|}\big{)}\times\texttt{sgn}(2|v_{ia}|-\\ &|v_{ib}|-|v_{ic}|)\times\Big{(}2\frac{\partial|v_{ia}|}{\partial x_{j\phi}}-\frac{\partial|v_{ib}|}{\partial x_{j\phi}}-\frac{\partial|v_{ic}|}{\partial x_{j\phi}}\Big{)}-\\ &\big{|}2|v_{ia}|-|v_{ib}|-|v_{ic}|\big{|}\times\Big{(}\sum_{i\phi}\frac{\partial{|v_{i\phi}|}}{\partial x_{j\phi}}\Big{)}\Big{\}}\end{aligned}$ (9)
@@
Similarly for the second case ($\big{|}|v_{ib}|-{v_{p}}_{i}^{\text{avg}}\big{|}>\big{|}|v_{ia}|-{v_{p}}_{i}^{\text{avg}}\big{|}~{}\&~{}\big{|}|v_{ic}|-{v_{p}}_{i}^{\text{avg}}\big{|}$), the $\text{PVUR}_{ij\phi}=$
$&\frac{\partial(\Xi^{\text{PVUR}})}{\partial x_{j\phi}}=\frac{1}{\big{(}\sum_{i\phi}{|v_{i\phi}|}\big{)}^{2}}\Big{\{}\big{(}\sum_{i\phi}{|v_{i\phi}|}\big{)}\times\texttt{sgn}(2|v_{ib}|-\\ &|v_{ia}|-|v_{ic}|)\times\Big{(}2\frac{\partial|v_{ib}|}{\partial x_{j\phi}}-\frac{\partial|v_{ia}|}{\partial x_{j\phi}}-\frac{\partial|v_{ic}|}{\partial x_{j\phi}}\Big{)}-\\ &\big{|}2|v_{ib}|-|v_{ia}|-|v_{ic}|\big{|}\times\Big{(}\sum_{i\phi}\frac{\partial{|v_{i\phi}|}}{\partial x_{j\phi}}\Big{)}\Big{\}},\end{aligned}$ (10)
and third case $\big{|}|v_{ic}|-{v_{p}}_{i}^{\text{avg}}\big{|}>\big{|}|v_{ia}|-{v_{p}}_{i}^{\text{avg}}\big{|}~{}\&~{}\big{|}|v_{ib}|-{v_{p}}_{i}^{\text{avg}}\big{|}$, it is defined as, $\text{PVUR}_{ij\phi}=$
$&\frac{\partial(\Xi^{\text{PVUR}})}{\partial x_{j\phi}}=\frac{1}{\big{(}\sum_{i\phi}{|v_{i\phi}|}\big{)}^{2}}\Big{\{}\big{(}\sum_{i\phi}{|v_{i\phi}|}\big{)}\times\\ &\texttt{sgn}{(2|v_{ic}|-|v_{ia}|-|v_{ib}|)}\times\Big{(}2\frac{\partial|v_{ic}|}{\partial x_{j\phi}}-\frac{\partial|v_{ia}|}{\partial x_{j\phi}}-\frac{\partial|v_{ib}|}{\partial x_{j\phi}}\Big{)}-\\ &\big{|}2|v_{ic}|-|v_{ia}|-|v_{ib}|\big{|}\times\Big{(}\sum_{i\phi}\frac{\partial{|v_{i\phi}|}}{\partial x_{j\phi}}\Big{)}\Big{\}}.\end{aligned}$ (11)
@@ III-C1 Definition
It is defined by NEMA [4] and used mostly by the drives and motor manufacturers. LVUR is defined in similar way as PVUR, but with the line voltages. The LVUR is defined as
@@
where, ${v_{l}}_{i}^{\text{avg}}=\frac{1}{3}(|v_{iab}|+|v_{ibc}|+|v_{ica}|)$. LVUR can be re-written into three-cases as LVUR =
$\begin{dcases}\frac{\big{|}2|v_{iab}|-|v_{ibc}|-|v_{ica}|\big{|}}{|v_{iab}|+|v_{ibc}|+|v_{ica}|}\\ \qquad\quad\big{|}|v_{iab}|-{v_{l}}_{i}^{\text{avg}}\big{|}>\big{|}|v_{ibc}|-{v_{l}}_{i}^{\text{avg}}\big{|}~{}\&~{}\big{|}|v_{ica}|-{v_{l}}_{i}^{\text{avg}}\big{|}\\ \frac{\big{|}2|v_{ibc}|-|v_{iab}|-|v_{ica}|\big{|}}{|v_{iab}|+|v_{ibc}|+|v_{ica}|}\\ \qquad\quad\big{|}|v_{ibc}|-{v_{l}}_{i}^{\text{avg}}\big{|}>\big{|}|v_{iab}|-{v_{l}}_{i}^{\text{avg}}\big{|}~{}\&~{}\big{|}|v_{ica}|-{v_{l}}_{i}^{\text{avg}}\big{|}\\ \frac{\big{|}2|v_{ica}|-|v_{iab}|-|v_{ibc}|\big{|}}{|v_{iab}|+|v_{ibc}|+|v_{ica}|}\\ \qquad\quad\big{|}|v_{ica}|-{v_{l}}_{i}^{\text{avg}}\big{|}>\big{|}|v_{iab}|-{v_{l}}_{i}^{\text{avg}}\big{|}~{}\&~{}\big{|}|v_{ibc}|-{v_{l}}_{i}^{\text{avg}}\big{|}.\\ \end{dcases}$ (13)

### v2

@@ Abstract
Power distribution networks, especially in North America, are often unbalanced due to the mix of single-, two-, and three-phase networks as well as due to the high penetration of single-phase devices at the distribution level, such as electric vehicle (EV) chargers and single-phase solar plants. However, the network operator must adhere to the voltage unbalance levels within the limits specified by IEEE, IEC, and NEMA standards for the safety of the equipment as well as the efficiency of the network operation. Existing works have proposed active and reactive power control in the network to minimize imbalances. However, these optimization problems are highly nonlinear and nonconvex due to the inherent nonlinearity of unbalanced metrics and power-flow equations. In this work, we propose a linearization approach of unbalance metrics such as voltage unbalance factors (VUF), phase voltage unbalance rate (PVUR), and line voltage unbalance rate (LVUR) using the first-order Taylor’s approximation. This linearization is then applied to the phase balancing control scheme; it is formulated as a feedback approach where the linearization is updated successively after the active/reactive control setpoint has been actuated and shows improvement in voltage imbalances. We demonstrate the application of the proposed scheme on a standard IEEE benchmark test case, demonstrating its effectiveness.
@@ I Introduction
The power distribution systems are progressively facing increased unbalanced power injections due to the rapid growth of single-phase solar plants and residential electric vehicle chargers. This change, along with the presence of inherent unbalancedness (due to a mix of single and two-phase feeders) of North American grids, gives rise to a higher level of voltage unbalance. An increased level of unbalances in distribution systems may trigger protection devices, increase network losses, and damage power equipments (e.g., distribution transformers). A network with high power unbalance results in poor utilization as one of the phases is overloaded compared to others, and this may necessitate system upgrades [1]. Therefore, the system operators must satisfy the standards regarding the unbalancedness to ensure safe and reliable operation of power distribution networks, defined by IEEE [2], IEC [3], and NEMA [4].
@@
In the existing literature, several methods are proposed to address the issue of voltage imbalance; they can be broadly categorized into three groups. The first method involves investing in new devices, such as static or dynamic VAR compensators, shunt capacitors, tap changers, voltage regulators, etc. [5], which can be costly and take time to plan and install them. The second approach is the phase switching or swapping strategy [6, 7, 8], which involves reassigning certain loads to different phases at a bus to balance loads and reduce voltage unbalance. There are also hybrid approaches that combine phase switching with the use of static VAR compensators [9]. Utilities often employ these methods manually, where maintenance crews travel to sites and physically switch loads. This process is costly, requiring extensive planning and potentially costing tens of thousands of dollars per phase swap [10]. These balancing maneuvers are typically performed periodically, as load characteristics vary with seasonal changes [11, 12, 13, 14, 15]. The third method uses active and reactive power control from existing or newly installed solar inverters and other distributed energy resources (DERs). Many studies suggest regulating the active and reactive powers from solar PV inverters [15, 16]. This strategy may require investments in the communication and situational awareness infrastructure for real-time control [17]. Most of these approaches are usually non-convex due to the non-linear and non-convex nature of the unbalance metrics as well as the power flow equations. Except there are some works that have a proposed successive linearization, such as in [18], convexification via linearized distflow, e.g., [19, 20], etc.
@@
In this work, we propose to linearize the voltage unbalance metrics by computing the sensitivities of the voltage unbalance factors with respect to the control variables, enabling to have a linear phase balancing formulation using the first-order Taylor’s approximation. These coefficients are derived for different unbalance metrics such as voltage unbalance factors (VUF), phase voltage unbalance rate (PVUR), and line voltage unbalance rate (LVUR). The work shows the accuracy of the linearization on an IEEE benchmark network, then also shows how it can be used to formulate a linearized phase balancing problem. In summary, the key contributions of the presented work are (i) analytical derivation of the sensitivities of the different voltage unbalance factors with respect to control variables such as active and reactive powers, (ii) formulating a linearized phase balancing scheme for the active and reactive power control per phase, minimizing the voltage unbalance metrics.
@@
The paper is organized as follows. Section II introduces the key notations and the voltage magnitude sensitivities. Section III derives the sensitivity coefficients of different voltage unbalance metrics using the voltage magnitude sensitivities. Section IV demonstrates an application of the proposed sensitivity coefficients. Section V presents numerical validation, and finally, Section VI concludes the work.
@@ II-A Notations
We consider a distribution network with network indices $i\in\mathcal{N}$, where the node indices are contained in set $\mathcal{N}=\{1,\dots,N\}$. We consider a three-phase system with phases denoted by $a,b$ and $c$. The symbols $v_{ia},v_{ib}$ and $v_{ic}$ denote the complex voltages for phase $\phi\in\Phi=\{a,b,c\}$ and the symbols $v_{iab}$, $v_{ibc}$ and $v_{ica}$ refer to line voltages. The symbols $p_{ia},p_{ib}$ and $p_{ic}$ and $q_{ia},q_{ib}$ and $q_{ic}$ denote the active and reactive powers, respectively. The symbols $|y|$ and $y^{*}$ denote the magnitude and complex conjugate of $y$, respectively.
@@ II-B Voltage Sensitivities
Here, $x$ can be active, reactive power, or other controllable variables such as tap changers. The same coefficient can be defined for voltage magnitudes and angles. The sensitivities of voltage magnitudes at bus $i$ and phase $\psi\in{\Phi}$ with respect to control variable at bus $j$, phase $\phi\in\Phi$ can be found as
$\frac{\partial{|v_{i\psi}|}}{\partial x_{j\phi}}=\frac{1}{|v_{i\psi}|}\Re\Big(v_{i\psi}^{*}\frac{\partial{v_{i\psi}}}{\partial x_{j\phi}}\Big).$ (2)
These sensitivity coefficients can either be computed by inverting the Jacobian of the power-flow solution using the Newton-Raphson approach or by an admittance matrix-based approach, which relies on solving a linear set of equations [21, 22, 23] or an impedance matrix-based approach as described in [24].
@@ III Voltage Unbalance Metrics and their Sensitivities
In a multi-phase electric distribution grid, the voltage unbalance is usually referred to the cases when the three-phase voltages are not symmetrical, which means that either voltage magnitudes are unequal or the voltage phase angles are not 120 apart or both of the above. This may happen due to inherent unbalancedness in the network, such as the presence of single- and two-phase circuits, untransposed circuits, or the presence of single-phase injections (load or generation). The degree of unbalance is quantified by several metrics defined by the international institutions such as IEEE[2], IEC [3] and NEMA [4]. These standards also prescribe the limits on each of these metrics for safe operation of the power distribution systems. In practice, the electric distribution system operators try their best to keep the system “balanced” within the limits defined by these standards.
In the following section, we introduce different voltage unbalance metrics and show how their sensitivity with respect to a generic control variable can be analytically expressed.
@@ III-A1 Definition
Here, $\alpha=e^{j\frac{2\pi}{3}}$ and symbols $v_{i0},v_{i+}$ and $v_{i-}$ denote the zero, positive, and negative sequence voltage, respectively, for $i$-th node. The IEC standard in [3] mandates keeping the VUF value within 0.02 per unit (or 2%) for the power distribution systems.
@@ III-B1 Definition
PVUR is defined by IEEE [2], and it uses the phase-to-ground voltages to express the imbalance metric instead of the sequence voltages, as was for the VUF. It is defined as the ratio of (i) the maximum deviation of the phase voltage magnitudes from their average and (ii) the average phase voltage magnitudes. It is $\Xi_{i}^{\text{PVUR}}=$
@@
[…]
@@ III-B2 PVUR Sensitivities (PVURS)
Based on the three cases defined in (8), the PVUR sensitivity for each of these cases can be computed by differentiating with respect to the control variables. Here, $\phi,\psi$ varies over the phase set $\Phi$ defined earlier.
[…]
$\begin{aligned} &\frac{\partial(\Xi_{i}^{\text{PVUR}})}{\partial x_{j\phi}}=\frac{1}{\big(\sum_{\psi}{|v_{i\psi}|}\big)^{2}}\Big\{\big(\sum_{\psi}{|v_{i\psi}|}\big)\times\texttt{sgn}(2|v_{ia}|-\\ &|v_{ib}|-|v_{ic}|)\times\Big(2\frac{\partial|v_{ia}|}{\partial x_{j\phi}}-\frac{\partial|v_{ib}|}{\partial x_{j\phi}}-\frac{\partial|v_{ic}|}{\partial x_{j\phi}}\Big)-\\ &\big|2|v_{ia}|-|v_{ib}|-|v_{ic}|\big|\times\Big(\sum_{\psi}\frac{\partial{|v_{i\psi}|}}{\partial x_{j\phi}}\Big)\Big\}\end{aligned}$ (9)
@@
[…]
$\begin{aligned} &\frac{\partial(\Xi^{\text{PVUR}})}{\partial x_{j\phi}}=\frac{1}{\big(\sum_{\psi}{|v_{i\psi}|}\big)^{2}}\Big\{\big(\sum_{\psi}{|v_{i\psi}|}\big)\times\texttt{sgn}(2|v_{ib}|-\\ &|v_{ia}|-|v_{ic}|)\times\Big(2\frac{\partial|v_{ib}|}{\partial x_{j\phi}}-\frac{\partial|v_{ia}|}{\partial x_{j\phi}}-\frac{\partial|v_{ic}|}{\partial x_{j\phi}}\Big)-\\ &\big|2|v_{ib}|-|v_{ia}|-|v_{ic}|\big|\times\Big(\sum_{\psi}\frac{\partial{|v_{i\psi}|}}{\partial x_{j\phi}}\Big)\Big\},\end{aligned}$ (10)
[…]
$\begin{aligned} &\frac{\partial(\Xi^{\text{PVUR}})}{\partial x_{j\phi}}=\frac{1}{\big(\sum_{\psi}{|v_{i\psi}|}\big)^{2}}\Big\{\big(\sum_{\psi}{|v_{i\psi}|}\big)\times\\ &\texttt{sgn}{(2|v_{ic}|-|v_{ia}|-|v_{ib}|)}\times\Big(2\frac{\partial|v_{ic}|}{\partial x_{j\phi}}-\frac{\partial|v_{ia}|}{\partial x_{j\phi}}-\frac{\partial|v_{ib}|}{\partial x_{j\phi}}\Big)-\\ &\big|2|v_{ic}|-|v_{ia}|-|v_{ib}|\big|\times\Big(\sum_{\psi}\frac{\partial{|v_{i\psi}|}}{\partial x_{j\phi}}\Big)\Big\}.\end{aligned}$ (11)
@@ III-C1 Definition
It is defined by NEMA [4] and used mostly by the drives and motor manufacturers. LVUR is defined in a similar way as PVUR, but with the line voltages. The LVUR is defined as
@@
where, ${v_{l}}_{i}^{\text{avg}}=\frac{1}{3}(|v_{iab}|+|v_{ibc}|+|v_{ica}|)$. LVUR can be rewritten into three cases as LVUR =
[…]